OpenStudy (anonymous):

Find the surface area of the sphere with the given dimension. Leave your answer in terms of . 18. radius of 70 m (1 point) 4,900 m2 19,600 m2 9,800 m2 14,700 m2

OpenStudy (anonymous):

Do you know the formula for the surface area of a sphere?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i think its B will explain it tho step by step

OpenStudy (anonymous):

You should memorize it. It's \[A=4 \pi r^2\] where r is the radius, of course.

OpenStudy (anonymous):

i got 61544 buts its not an answer choice

OpenStudy (anonymous):

Remember that it asked you to leave the answer in terms of pi, which means that you don't multiply an approximation of pi, so your answer should look like _____*pi.

OpenStudy (anonymous):

uhhh yeah idk how to do any of this

OpenStudy (anonymous):

Look at the formula: A = 4 * pi * r^2 You have the fact that r=70, so plug it in. A = 4 * pi * 70^2 A = 4 * pi * 4900 You want your answer to look like A = _____*pi m^2.

OpenStudy (anonymous):

so its 4900

OpenStudy (anonymous):

9. A spherical balloon has a circumference of 21 cm. What is the approximate surface area of the balloon to the nearest square centimeter? (1 point) 1,385 cm2 561 cm2 346 cm2 140 cm2

OpenStudy (anonymous):

Not quite. A = 4 * pi * 4900 It needs to simplify so that there's only one string of numbers, with pi after it.

OpenStudy (anonymous):

im lost i really dont get this so its not 4900

OpenStudy (anonymous):

It's not 4900. Look at what you have: A = 4 * pi * 4900 If you reorder it, that's A = 4 * 4900 * pi You can simplify it so that it's in "terms of pi" from here. Remember that it should look like: A = ____ * pi

OpenStudy (anonymous):

so its 4900* pi?

OpenStudy (anonymous):

No, you're still missing something. A = 4*4900 * pi

OpenStudy (anonymous):

i got 19,600 ?

OpenStudy (anonymous):

Yes! 19600pi m^2 is the answer.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

As for your second question, let's look at what we have: C = 21 cm A = 4*pi*r^2 We can't plug the circumference into this formula, though. You have to find the radius now, from the circumference formula: C = 2*pi*r That means: 21 = 2*pi*r Now we need to solve for the radius, r.

OpenStudy (anonymous):

how to find the radius

OpenStudy (anonymous):

You have 21=2*pi*r Use algebra to isolate r.

OpenStudy (anonymous):

im so confused will u pls explain.... i am dyslexic

OpenStudy (anonymous):

21 = 2 * pi * r Remember that in algebra, we do the opposite function to the opposite function to move things to the other side. This is "isolating the variable." For example, when on one side we're multiplying, we can divide both sides to move it to the other side. 21 = 2 * pi * r To get rid of the 2 we're multiplying on the right, we divide both sides by 2, and we get: 21/2 = 2/2 * pi * r This simplifies down to: 10.5 = pi * r To get rid of the pi we're multiplying on the right, we divide both sides by pi, and we get: 10.5/pi = pi/pi * r This simplifies down to: 3.342 = r r = 3.342 That's the radius measure.

OpenStudy (anonymous):

To find the surface area, we use the same process we used in the first question, except it wants us to approximate the answer (that means it'll have decimals, __.__ instead of ___*pi) now. A = 4 * pi * r^2 We have the fact that r = 3.342 Plug it in to get: A = 4 * pi * 3.342^2 A = 4 * pi * 11.169 You can multiply all of those now, since we want the approximation.

OpenStudy (anonymous):

140

OpenStudy (anonymous):

Yup! :)

OpenStudy (anonymous):

21. The volume of a sphere is 4,000 m3. What is the surface area of the sphere to the nearest square meter? (1 point) 2,614 m2 3,836 m2 1,307 m2 794 m2

OpenStudy (anonymous):

im badd at this

OpenStudy (anonymous):

This is another one where we'll need to solve for r. We have: V = 4000 And we have the formulas: A = 4 * pi * r^2 V = (4/3) * pi * r^3 Using this knowledge, we have: 4000 = (4/3) * pi * r^3 This is going to need us to use algebra again. First, let's get rid of the 4/3 on the right side. We'll divide both sides by it. 4000 / (4/3) = (4/3)/(4/3) * pi * r^3 This simplifies down to 3000 = pi * r^3 Again, we can get rid of the pi by dividing both sides by it. 3000/pi = pi/pi * r^3 This, now, simplifies to: 954.93 = r^3 On the right hand side, we're "cubing" r, or ^3. To undo this, like we undo a "square" or ^2, we'll take a root of it. Undoing ^2 means that we take the square root. Likewise, undoing ^3 means that we take the cube root. \[\sqrt[3]{954.93}=\sqrt[3]{r^3}\]This simplifies down to 9.847 = r r = 9.847

OpenStudy (anonymous):

Again, it's asking for an approximation, rounded to the nearest m^2. Plug it in to the formula: A = 4 * pi * r^2 A = 4 * pi * (9.847)^2 That should simplify nicely.

OpenStudy (anonymous):

i didnt get any of the answers ugggghhh

OpenStudy (anonymous):

i got 1217?

OpenStudy (anonymous):

I got that too. I googled it and there's nothing wrong with my math. I don't see why that's not an answer. Try the closest one (1307), and report it to your teacher I guess.

OpenStudy (anonymous):

i have a lot of different answers than my teacher ima post them and will u pls check them?

OpenStudy (anonymous):

alright I guess

OpenStudy (anonymous):

9. What are the lateral area and the surface area of the cone shown below? Round the answers to the nearest tenth. The figure is not drawn to scale. a cone (1 point) LA = 615.8 ft2; SA = 1,803.3 ft2 LA = 615.8 ft2; SA = 3,166.7 ft2 LA = 1,187.5 ft2; SA = 1,803.3 ft2 LA = 1,187.5 ft2; SA = 3,166.7 ft2

OpenStudy (anonymous):

r is 14 and h is 27 i got SA = 1952.3 and La = 1337.7

OpenStudy (anonymous):

Those answers are both correct.

OpenStudy (anonymous):

they arent on there ima call my teacher thank u so much

OpenStudy (anonymous):

For questions 11 and 12, find the volume of the cylinder in terms of . 11. a cylinder (1 point) 43.35 m3 51 m3 56.78 m3 73.695 m3 12. h = 10 and r = 5 (1 point) 625 in.3 500 in.3 125 in.3 250 in.3

OpenStudy (anonymous):

for 11 the r is 1.7 and the h is 15 and i got 136.19 and idk 12

OpenStudy (anonymous):

12 i got 785

OpenStudy (anonymous):

11 and 12 are different, they want answers in terms of pi again. 11 simplifies to 136.19, but it needs to be in terms of pi, so it's 43.35pi m^3. 12 simplifies to 785 just the same, but it needs to be in terms of pi, so it's 250pi in^3.

OpenStudy (anonymous):

21 and 9 you should ask about though. If we're wrong, your teacher will hopefully explain the answers.

OpenStudy (anonymous):

13. What is the height of the cylinder? The figure is not drawn to scale. a cylinder (1 point) 14.7 in 6.6 in 2.1 in 4.2 in r is 7

OpenStudy (anonymous):

v is 323.5

OpenStudy (anonymous):

Here's the formula: V = h*pi*r^2 I'm gonna run through the algebra: 323.5 = h*pi*(7)^2 323.5 = h*pi*49 323.5/49 = h*pi*49/49 6.6 = h*pi 6.6/pi = h*pi/pi 2.1 = h

OpenStudy (anonymous):

thnks i got 2.1 i jus wanted to check

OpenStudy (anonymous):

14. What is the volume of the composite space figure to the nearest whole number? The figure is not drawn to scale. a composite space figure (1 point) 216 cm3 90 cm3 378 cm3 306 cm3

OpenStudy (anonymous):

Add the volumes of those two rectangular prisms (the boxy shapes). It looks like the one on the left has the dimensions 4*9*6. The one on the right has 5*3*6. V = (4*9*6)+(5*3*6)

OpenStudy (anonymous):

306?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

For questions 22 and 23, determine whether the two figures are similar. If so, give the similarity ratio of the smaller figure to the larger figure. yes; 1:1.2 yes; 1:1.4 yes; 1:2.4 no

OpenStudy (anonymous):

i cant thank u enough

OpenStudy (anonymous):

25. If the scale factor of two similar solids is 3:16, what is the ratio of their corresponding areas and volumes? (1 point) 3:256 and 3:4,096 27:4,096 and 9:256 6:32 and 9:48 9:256 and 27:4,096 26. The volumes of two similar solids are 1,331 m3 and 216 m3. The surface area of the larger solid is 484 m2. What is the surface area of the smaller solid? (1 point) 864 m2 288 m2 144 m2 68 m2

OpenStudy (anonymous):

121:1,936 1:64 4:1 1:4 25. If the scale factor of two similar solids is 3:16, what is the ratio of their corresponding areas and volumes? (1 point) 3:256 and 3:4,096 27:4,096 and 9:256 6:32 and 9:48 9:256 and 27:4,096 26. The volumes of two similar solids are 1,331 m3 and 216 m3. The surface area of the larger solid is 484 m2. What is the surface area of the smaller solid? (1 point) 864 m2 288 m2 144 m2 68 m2

OpenStudy (anonymous):

24. What is the scale factor of a cube with volume 1,331 ft.3 to a cube with volume 85,184 ft.3? (1 point) 121:1,936 1:64 4:1 1:4

OpenStudy (anonymous):

looks like 22 is no

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

sorry, I have to go now

OpenStudy (anonymous):

ok:(

OpenStudy (anonymous):

thanks